jabez177 (jabez177):

The product of 3 and a number is 12 more than the sum of the number and 26. What is the number? I really need the asnwer because i have to answer this and then leave to go to my dental appointment

jabez177 (jabez177):

The product of 3 and a number is 12 more than the sum of the number and 26. What is the number?

OpenStudy (anonymous):

4

jabez177 (jabez177):

Okay thank you so much @Livvie_Marie_Belmont

OpenStudy (anonymous):

your welcome @jabez177

OpenStudy (mathstudent55):

Let x be the number. "The product of 3 and a number" means 3x "is" means = "12 more than the sum of the number and 26" means (x + 26) + 12 Put it all together to get: 3x = (x + 26) + 12 3x = x + 26 + 12 3x = x + 38 2x = 38 x = 19 The answer is "the number is 19."